x1 6 11 15 18 20
y1 6 8 12 20 30

Compute b1 and b0 (to 3 decimals). Enter negative values as negative numbers.

b1

b0

Complete the estimated regression equation (to 3 decimals). Enter negative values as negative numbers.

y^ = ____ + _____ x

x1

x2

x3

x4

x5

Compute the standardized residuals. Enter negative values as negative numbers.

standardized residual
1   (to 4 decimal places)
2   (to 4 decimal places)
3   (to 4 decimal places)
4   (to 4 decimal places)
5   (to 4 decimal places)
Answer #1
Line of Regression Y on X i.e Y = bo + b1 X
X Y (Xi - Mean)^2 (Yi - Mean)^2 (Xi-Mean)*(Yi-Mean)
6 6 64 84.64 73.6
11 8 9 51.84 21.6
15 12 1 10.24 -3.2
18 20 16 23.04 19.2
20 30 36 219.04 88.8

calculation procedure for regression
mean of X = sum ( X / n ) = 14
mean of Y = sum ( Y / n ) = 15.2
sum ( (Xi - Mean)^2 ) = 126
sum ( (Yi - Mean)^2 ) = 388.8
sum ( (Xi-Mean)*(Yi-Mean) ) = 200
b1 = sum ( (Xi-Mean)*(Yi-Mean) ) / sum ( (Xi - Mean)^2 )
= 200 / 126
= 1.587
bo = sum ( Y / n ) - b1 * sum ( X / n )
bo = 15.2 - 1.587*14 = -7.022
value of regression equation is, Y = bo + b1 X
Y'=-7.022+1.587* X          
------------------------------------------------------------------------------
b1 = 1.587 ;
bo = -7.022
regression equation = Y'=-7.022+1.587* X      

RESIDUAL CHART

Observation Y Predicted Residual
1 6.0 2.5 3.5
2 8.0 10.4 -2.4
3 12.0 16.8 -4.8
4 20.0 21.5 -1.5
5 30.0 24.7 5.3